# Pulse Shaping Filter Design for Integrated Sensing & Communication with Zak-OTFS

### v1

@@ IV-B Pulse Shaping Filter Visualization
Fig. 1 visualizes the pulse shaping filter magnitudes and their energy spectral densities (ESD). All pulse shapes have $99.99\%$ energy concentration within $f\in[\nicefrac{{-B}}{{2}},\nicefrac{{B}}{{2}}]$ and $t\in[\nicefrac{{-T}}{{2}},\nicefrac{{T}}{{2}}]$. We observe that the proposed IOTA pulse shapes are more localized than the Gaussian-sinc filter [8]. The IOTA PSWF pulse is also more localized than the IOTA Gaussian filter due to the optimality of the PSWF pulse in localization [14, 15].
IV-C Performance with Perfect I/O Relation
@@
Fig. 2 plots the bit error rate (BER) for uncoded $4$-QAM (quadrature amplitude modulation) transmissions, assuming perfect I/O relation knowledge at the receiver. The proposed IOTA pulses offer ideal BER performance similar to the sinc filter due to the orthogonality of the pulse. With perfect I/O relation knowledge, we observe no significant performance difference between the sinc, the Gaussian-sinc [8] and the proposed IOTA pulses due to their near-orthogonality. However, the Gaussian filter has poor BER due to its non-orthogonality.
IV-D Performance with Estimated I/O Relation
Fig. 3 plots the uncoded $4$-QAM BER and normalized mean squared error (NMSE) for I/O relation estimation using a separate pilot frame with pilot signal-to-noise ratio (SNR) equal to the data SNR.
The non-localized behavior of the sinc filter results in poor I/O relation estimation, degrading the BER and NMSE. Due to its near-ideal localization, the NMSE corresponding to the Gaussian filter reduces exponentially with the SNR; however, its non-orthogonality results in poor BER performance. The Gaussian-sinc filter [8] achieves significantly better NMSE and BER compared to the sinc filter. However, since the filter is not strictly orthogonal, the NMSE does not reduce exponentially as a function of SNR and plateaus at high SNRs – resulting in sub-optimal BER performance. In contrast to the above pulse shapes, the simultaneous localization and orthogonality of the proposed IOTA pulses results in NMSE close to that of the Gaussian filter (with exponential reduction with SNR) and significant improvement in BER compared to the Gaussian-sinc filter. Moreover, the better localization of the IOTA PSWF filter over the IOTA Gaussian filter (see Fig. 1) results in better BER and NMSE performance. We note that the proposed IOTA filters achieve ideal NMSE and BER performance without any time or bandwidth expansion, unlike the RRC filter with $1.6\times$ larger time and bandwidth.
@@ V Conclusion
In this paper, we designed optimal delay-Doppler pulse shaping filters for integrated sensing & communication via the Isotropic Orthogonal Transform Algorithm. In contrast to prior approaches that do not meet all three objectives for pulse shaping filter design – localization, orthogonality on the information lattice, and bandwidth/time limitation – the proposed filters achieve all three objectives simultaneously; thus offering ideal integrated sensing & communication performance. Future work will explore hardware demonstrations of the proposed designs and information theoretically quantify the sensing-communication trade-off in Zak-OTFS.

### v2

@@ IV-B Pulse Shaping Filter Visualization
Fig. 1 visualizes the pulse shaping filter magnitudes and their energy spectral densities (ESD), excluding the RRC pulse. All pulse shapes have $99.99\%$ energy concentration within $f\in[\nicefrac{{-B}}{{2}},\nicefrac{{B}}{{2}}]$ and $t\in[\nicefrac{{-T}}{{2}},\nicefrac{{T}}{{2}}]$. We observe that the proposed IOTA pulse shapes are more localized than the Gaussian-sinc filter [2]. The IOTA PSWF pulse is also more localized than the IOTA Gaussian filter due to the optimality of the PSWF pulse in localization [20, 9].
IV-C I/O Relation Estimation (Sensing) Performance
Fig. 2 plots the normalized mean squared error (NMSE) for estimating the DD effective channel $\mathbf{h}_{\mathrm{eff}}[k,l]$ (sensing) via (II) using a separate pilot frame with pilot signal-to-noise ratio (SNR) equal to the data SNR. The NMSE is defined as $\text{NMSE}=\frac{\sum_{k,l}|\widehat{\mathbf{h}}_{\mathrm{eff}}[k,l]-\mathbf{h}_{\mathrm{eff}}[k,l]|^{2}}{\sum_{k,l}|\mathbf{h}_{\mathrm{eff}}[k,l]|^{2}}$ for an estimate $\widehat{\mathbf{h}}_{\mathrm{eff}}[k,l]$ of the DD effective channel. We observe that maximum localization of the Gaussian and RRC pulses (due to $1.6\times$ time and bandwidth expansion) results in the best NMSE performance (exponential reduction with SNR), whereas the non-localized behavior of the sinc pulse results in the worst NMSE performance (plateaus at moderate SNRs). The Gaussian-sinc filter [2] achieves significantly better NMSE compared to the sinc filter; however, since it is not strictly orthogonal, the NMSE plateaus at high SNRs. On the other hand, the improved localization of the proposed IOTA pulses over the Gaussian-sinc pulse (see Fig. 1) results in improved NMSE performance close to the Gaussian filter, with exponential reduction vs SNR.
IV-D Data Detection Performance
@@
Figs. 3(3(a))-(3(b)) plot the bit error rate (BER) for uncoded $4$-QAM (quadrature amplitude modulation) transmissions for perfect and estimated I/O relation knowledge at the receiver respectively. I/O relation estimation is performed similar to Section IV-C, i.e., using a separate pilot frame with pilot SNR equal to the data SNR. Note that the Gaussian filter has poor BER due to its non-orthogonality, and the poor NMSE for I/O relation with the sinc filter (Fig. 2) results in poor BER with estimated I/O relation. Due to their orthogonality, localized behavior, and improved NMSE performance from Fig. 2, the proposed IOTA pulses offer ideal BER performance with both perfect and estimated I/O relation, with improved performance due to lower NMSE over the Gaussian-sinc filter from [2].
IV-E Spectral Efficiency Comparison of Various Pulse Shapes
Figure 4: Spectral efficiency of various pulse shaping filters for uncoded $4$-QAM transmission with I/O relation estimation.
Fig. 4 compares the spectral efficiency (SE) achieved by different pulse shapes for uncoded $4$-QAM transmission with estimated I/O relation. Due to $1.6\times$ time and bandwidth expansion, the RRC filter achieves the lowest SE, with a maximum value of $\nicefrac{{\log_{2}{4}}}{{1.6^{2}}}\approx 0.78$ bits/s/Hz. Due to its non-orthogonality and poor BER performance in Fig. 3, the Gaussian filter achieves the maximum possible SE of $2$ bits/s/Hz only at $30$ dB SNR. Due to poor I/O relation estimation and BER saturation at $10^{-2}$ in Fig. 3(3(b)), the sinc filter achieves a maximum SE of $0.99\times 2=1.98$ bits/s/Hz. The proposed IOTA filters achieve the best SE performance, with improvements of about $1\%$ over the Gaussian-sinc SE at $10$ dB SNR.
@@ V Conclusion
In this paper, we applied the Isotropic Orthogonal Transform Algorithm to the design of delay-Doppler pulse shaping filters for integrated sensing & communication. Unlike previous approaches, the proposed filters are localized, achieve full spectral efficiency and do not suffer from inter-symbol interference. Future work will investigate the interplay between pulse shaping and coding, and explore extensions to multi-user MIMO and applications such as sidelink-based sensing.